Hi! Can you please help me solve this problem?
I need to find the area of a right angled triangle involving surds. The points are A, B & C. The height of this triangle is √5+3/2 and the length is 2. How would I work this out? I know I have to use this formula: 1/2 x base x height,but the surds are confusing me.
Thank you so much in advanced!

Answer by richwmiller(17219)

1/2 base =1/2*2=1
so the area is the surd!